Wardell Company purchased a minicomputer on January 1, 2016, at a cost of $49,000. The computer was depreciated using the straight-line method over an estimated five-year life with an estimated residual value of $7,000. On January 1, 2018, the estimate of useful life was changed to a total of 10 years, and the estimate of residual value was changed to $400.

Required:
1. Prepare the year-end journal entry for depreciation in 2018. No depreciation was recorded during the year.
2. Prepare the year-end journal entry for depreciation in 2018. Assume that the company uses the sum-of-the-years' -digits method instead of the straight-line method.
Answer #1
CALCULATION OF THE DEPRECIATION AS PER STRAIGHT LINE METHOD
Purchase Cost of Mini Computer $                      49,000
Less: Residual Value $                        7,000
Value for the depreciaiton $                      42,000
Divide by useful life = "/" 5 Years  
Depreciation per year ($ 42,000 / 5 Years) $                        8,400
CACLULATION OF THE BOOK VALUE AS ON JANUARY 01, 2018 AND DEPRECIATION AS PER SLM
Purchase Price = $                      49,000
Less: Depreciation for the year 2016 $                        8,400
Less: Depreciation for the year 2017 $                        8,400
Book Value at the year end december 31, 2017= $                      32,200
Opening balance as on Jan, 01 2018 = $                      32,200
Less: Salvage Value = $                            400
Value for the depreciation $                      31,800
Balance life of the assets (10 Years - 2 Years) 8 years
Depreciation Per year ($ 31,800 / 8 Years ) $                        3,975
So depreciation for the year 2018 as per Striaght Line method = $                        3,975
CALCULATION OF THE DEPRECIATION AS PER SUM OF DIGIT METHODS
Opening balance as on Jan, 01 2018 = $                      32,200
Less: Salvage Value = $                            400
Value for the depreciation $                      31,800
Sum of Years Digit = 1+ 2 + 3 + 4 +5 + 6+ 7 + 8 = 36
So, Depreciation for the year 2018 = $ 31,800 X 8 / 36 = $                  7,066.67
Journal Entries
Date Account Title and explanation Debit Credit
Answer = 1)
2018 Depreciation Expenses $                        3,975
     To Accumulated Depreciation - Microcomputer $                       3,975
Answer = 2)
2018 Depreciation Expenses $                  7,066.67
     To Accumulated Depreciation - Microcomputer $                 7,066.67